Light takes 4.5 years to reach us from the nearest star. If the velocity of light is 3 ×10^8 ms-1, how far away is the star? A rocket first outside the Earth’s atmosphere in space accelerates at 30ms-2 for a week (7 days). How long will it take for the rocket to reach the star? (Ignore any gravitational attraction).


1
Expert's answer
2021-04-18T19:34:05-0400

i) v = 3"\times"108 "\frac{m}{s}"= 9.47"\times"1012 "\frac{km}{year}"

l = 9.467"\times"1012"\times"4.5 = 4.26"\times"1013 km


ii) v = 30"\times"(7"\times"24"\times"3600) = 1.8144"\times"107 "\frac{m}{s}" = 18144 "\frac{km}{s}"

t = "\frac{4.26\times10^{13}}{18144}" = 2.35"\times"109 years